The binomial theorem states that (a + b)n = Σ (prbn_k. (a) Use the binomial theorem to show that 2k-0 W = 2n. (Hint, 2n= (1 + 1)n.) (b) Expand (a2 + b)4.
Answer #1

According to Binomial theorem,

(a + b)n = nC0 an b0 + nC1 an-1 b1 + nC2 an-2 b2​ + ...... nCn a0 bn

Substituting a = 1 and b = 1 in the above theorem

=> (1 + 1)n = nC0 1n 10 + nC1 1n-1 11 + nC2 1n-2 12​ + ...... nCn 10 1n

=> 2n = nC0 * 1 + nC1 * 1 + nC2 * 1 + ...... nCn * 1

=> 2n = nC0 + nC1 + nC2 + ...... nCn

=> gif.latex?%5Csum_%7Bk%3D0%7D%5E%7Bn%7D%2 = 2n.

(b) (a2 + b)4

= 4C0 (a2)4 b0 + 4C1 (a2)3 b1 + 4C2 (a2)2 b2 + 4C3 (a2)1 b3 + 4C4 (a2)0 b4

= 1 * a8 * 1 + 4 * a6 * b + 6 * a4 * b2 + 4 * a2 * b3 + 1 * 1 * b4

= a8 + 4a6b + 6a4b2 + 4a2b3 + b4.